OpenStudy (anonymous):

Answer to the system by elimination -x - 7y=14. -4x - 14y=28

OpenStudy (zehanz):

First make a choice: do you want to eliminate x or y?

OpenStudy (anonymous):

(-x-7y=14)*-2 This now equals 2x+14y=-28 2x+14y=-28 -4x-14y=28 (subtract bottom equation from top) now you have -2x=0 x=0 Plug in x in an equation -4(0)-14y=28 -14y=28 y=-2 (0, -2)